Calculate the number of moles in 112.7 g of each of the following compounds.
SO2
Na2SO4
BaSO4
KAI(SO4)2
1
Expert's answer
2020-02-20T07:45:05-0500

Number of moles"(N)=\frac{M_g}{M_a}" where "M_g" is the given mass of compound and "M_a" is the molar mass of compound.

In the given question,"M_g=112.7 \ g" same for all compounds.

"SO_2:"

"M_a=32+2\times 16=64\ g"

"N=\frac{112.7}{64}=1.76"

"Na_2SO_4:"

"M_a=23\times 2+32+16\times 4=142\ g"

"N=\frac{112.7}{142}=0.794"

"BaSO_4:"

"M_a=137+32+16\times 4=233\ g"

"N=\frac{112.7}{233}=0.484"

"KAl(SO_4)_2:"

"M_a=39+27+2\times(32+16\times 4)=258\ g"

"N=\frac{112.7}{258}=0.437"